Roll over each name to read the description. Decide which of Hofstede’s cultural values is represented. Then drag the name to the appropriate area.

Power Distance

Individualism/ Collectivism

Uncertainty Avoidance

Masculinity/ Femininity

Confucian Dynamism

Elena- Elena has worked at the same company for more than 15 years. She parks same parking in the same parking spot every day and values routine and Predictability.

Ming - Ming is intellectually gifted and from a very powerful family in China. Due to power and wealth of her family, she is expected to be extremely successful in the workplace

Jiro -Jiro is extremely driven and assertive. His Goal is to become the most powerful man in the company.

Bill – Bill has a never give up attitude. He has a high degree of respect for tradition and will go to great lengths to protect his reputation.

Ben- Ben has very loose ties with his family and enjoys wearing provocative T-shirts to express a sense of freedom and uniqueness. He shows no interest in joining the family business.
Answer #1

Individualism/ Collectivism: It is an index which denotes the degree of associativity of people with the society. Individualism refers to giving importance to oneself whereas collectivism denotes a strong integration in the members of a society.

In this case Ben has very loose ties with family. He has no interest in family business. This is an example of individualism vs. collectivism.

Masculinity/Femininity: It is the preference of people especially males for the achievement and heroism in the society. In highly masculine society, no doubt women are more competitive and emphatic but the male people enjoy a higher status.

Thus, Jiro's case comes under Masculinity/Femininity.

Power Distance index is the degree to which the less powerful members believe in the unequal power distribution i.e. they believe that the distribution of power is not justified. Power distance is related to the acceptance level of the status and power of different people in the culture.

In this case, Ming is expecting to be successful at her workplace using the power and wealth of her family. Thus, in this case she can be considered under power distance.

Confucian Dynamism relates the past connections to predict current and future position of a person. It is also called as Long-term orientation vs. short-term orientation.

In this case Bill's high degree of respect for his culture and his tendency to protect his reputation shows that he wants to be in long term relationship with the concerned using his abilities. Thus, he can be considered under Confucian dynamism.

Uncertainty avoidance index states a society's tolerance for different opinions.

A high UA index means a strict adherence to the stiff guidelines, rules, and laws. Such a society usually rely on absolute Truth, or their own beliefs that are perceived to be truth. Whereas, a lower UA index depicts the acceptance of different opinions. In a low UA index society people are free to do whatever they wish to do.

In the case where Elena working for a long period of 15 years for the same company and using same parking spot, values routine and predictability shows that she adheres to formal rules.

Thus, Elena's case can be classified as uncertainty avoidance case.

$$ \begin{array}{|c|c|c|c|c|} \hline \text { Power Distance } & \begin{array}{c} \text { Individualism/ } \\ \text { Collectivism } \end{array} & \begin{array}{c} \text { Uncertainty } \\ \text { Avoidance } \end{array} & \begin{array}{c} \text { Masculinity/ } \\ \text { Femininity } \end{array} & \begin{array}{c} \text { Confucian } \\ \text { Dynamism } \end{array} \\ \hline \text { Ming } & \text { Ben } & \text { Elena } & \text { Jiro } & \text { Bill } \\ \hline \end{array} $$